OpenStudy (anonymous):

According to our rules for the order of operations, is this true or false? 6 - 5(8) = 8

myininaya (myininaya):

6-5(8) = 6-40 = -34

OpenStudy (anonymous):

false

OpenStudy (amistre64):

..... so much to say, so little time to say it lol